A cube with sides measuring 5cm is place in an electric field with magnitude 3x10^3 N/C.

a. Calculate the flux in each side of the six face of the cube.

b. Calculate the total flux of the cube


1
Expert's answer
2022-03-07T12:45:05-0500

(A)

Electric flux

"\phi=E.A=EAcos\theta"

"\phi=E\times a^2 cos0°"

"\phi=3\times10^3\times25\times10^{-4}=7.5Nm^2/C"

(B)

Net flux


"\phi_{net}=\phi_1+\phi_2+\phi_3+\phi_4+\phi_5+\phi_6"

"\phi_{net}=0Nm^2/C"